Plus Group has equipment with an original cost of Tic 50,00,000 and related accumulated depreciation Tk 10,00,000 on December 31, 2017. The fair value of the equipment at December 31, 2017, is lk 65,00,000. The equipment has a useful life of 5 years remaining after December 31, 2017, with having residual value of Tic 35,000. Plus uses the straight-line method of depreciation. Prepare the err to (i) record the revaluation of the equipment on December 31, 2017, and (ii) record depreciation on the equipment at December 31, 2018.


1
Expert's answer
2021-05-31T11:07:43-0400

i)


"65,000,000-(50,000,000+10,000,000)=5,000,000"

ii)


"\frac{65,000,000-35,000}{5}=12,993,000"